# Human foraging strategies flexibly adapt to resource distribution and time constraints

### v2

@@ 1 Introduction
Foraging is a nearly universal behaviour in the animal kingdom and essential for survival of most species (Mobbs et al.,, 2018; Hayden,, 2018; Hills,, 2006). Foraging problems encompass not just the search for survival-related resources (e.g., foraging for food) but also the search for information in the environment (e.g. in internet search (Pirolli,, 2007)) or in memory (e.g., “mental search” (Hills et al.,, 2015; Wilke et al.,, 2009; Todd and Hills,, 2020)).
In natural foraging problems, such as when animals navigate the environment to search for food (or other valuable resources), choice offers are presented serially and often found in clumps or patches, which gradually exhaust. This makes foraging scenarios conceptually different from classical economic tasks used to study value-based decisions, in which choice offers are typically presented simultaneously. Given their peculiarities, foraging choices are often conceptualized in terms of stay or leave decisions; namely, deciding whether to accept the available alternative (e.g., continue foraging in the same patch, in so-called patch-leaving problems) or reject it, in the hope that a better one would be found (e.g., move to another patch) (Charnov,, 1976; Hayden,, 2018; Hayden and Moreno-Bote,, 2018). Stay-or-leave problems are extensively studied, both in humans and animals (e.g., mammals (Blanchard and Hayden,, 2015; Kane et al.,, 2022; Hayden et al.,, 2011; Gancarz et al.,, 2023; Shahidi et al.,, 2024), birds (Stephens and Dunlap,, 2009; Chandel et al.,, 2021)). Similarly, the trade-offs between exploration and exploitation that arise during foraging and related tasks have been widely studied in humans and other animals (Cohen et al.,, 2007; Mehlhorn et al.,, 2015; Addicott et al.,, 2017; Daw et al.,, 2006; Verschure et al.,, 2014; Maffei et al.,, 2015; Parr et al.,, 2022; Pezzulo et al.,, 2018; Friston et al.,, 2017)).
A formal solution to patch-leaving and other foraging problems is the Marginal Value Theorem of optimal foraging theory (Charnov,, 1976), according to which the most profitable moment to leave a patch (“giving-up time”) is when the reward in the current patch is less than the average reward rate of the environment (Constantino and Daw, 2015a, ). Neurally, the decision to stay or to leave is influenced by the time-linked reward representations at the level of the dACC (Wittmann et al.,, 2016). Furthermore, a recent PET study has demonstrated the role of dopamine, essential for motivational states that drive human decision-making dynamics (Le Heron et al.,, 2019), in regulating behavioral adaptations to changes in the reward environment by weighing reward benefits against temporal costs (Ianni et al.,, 2023).
@@
A series of human studies tested how environmental constraints – the most important being resource distribution – affect individuals’ patch-leaving strategies. Using a multiple-target visual search task, Cain and colleagues (Cain et al.,, 2012) demonstrated that foragers adapt to environmental statistics in a way that is consistent with, although not identical to, the optimal formulation: participants search more when expecting a greater amount of targets and modify their expectations in real time by considering the target distributions across trials. In a different experiment, a virtual patch-foraging task was used to assess the influence of environmental richness on foraging decisions (Constantino and Daw, 2015b, ). Results have shown that although participant’s behavior was in agreement with the optimal policy there was an “overharvesting” tendency, namely, a tendency to harvest a patch longer than it would be needed for reward maximization; see also (Hutchinson et al.,, 2008). Overharvesting has been explained as a strategy for reducing the cost of behavioral variability (Cash-Padgett and Hayden,, 2020), as a result of optimal learning and adaptability (Harhen and Bornstein,, 2023), or as a phenomenon occurring as a consequence of temporal cognition (Kendall and Wikenheiser,, 2022). Indeed, time definitely matters when making patch-leaving decisions; as the MVT states, longer travel times lead to longer patch residence times, which lowers the forager’s opportunity cost of time as well as the environment’s overall rate of reward (Bustamante et al.,, 2023).
@@
Like resources are limited in the environment, time is also a finite resource. Animals allocate different time foraging, depending on external constraints and they flexibly adjust their foraging behavior to time available (Krebs,, 1980; Swennen et al.,, 1989; Ydenberg et al.,, 1994; Ydenberg and Hurd,, 1998). For example, they tend to work harder when time available for foraging is more limited (Lewis et al.,, 2004). Time pressure also implies that decisions have to be made faster. A recent study employed a time bandit task to examine the effects of time pressure on exploration and decision making, found that in limited time conditions, participants earned less reward, were less sensitive to reward values in repeat choice behavior, were less likely to choose options that were associated with higher uncertainty, and explored less in uncertain scenarios. These findings can be explained by considering that under time pressure, participants tend to select simpler, lower-cost policies (Wu et al.,, 2022).
@@
In sum, the above studies contributed to shed light on how two environmental constraints – namely, resource distribution and time availability – affect human patch-leaving strategies. However, these two constraints have been manipulated in isolation, leaving open the question of how they might interact and whether participants are sensitive to both factors. Finally, most human foraging studies used simple paradigms, in which rewards are paired to buttons that are simultaneously present on the screen. It is unclear how well their results generalize to foraging scenarios that resemble more closely animal patch-living problems, in which rewards are presented serially, visibility is partial and which engage sophisticated cognitive processes, such as those required to learn and remember reward locations, orienting and navigating in space, etc. (Mobbs et al.,, 2018; Hayden,, 2018; Yoo et al.,, 2021; Cisek and Pastor-Bernier,, 2014).Crucially, these more challenging scenarios present various sources of uncertainty about resource locations and distributions – for example, how many reward locations can be reached within limited time, where rewards are located in space and how to navigate more efficiently to reach them – and hence multiple occasions for learning. Therefore, they would allow studying whether participants improve foraging performance over time and whether they match (or approximate well) optimal solutions from foraging theory (Allen et al.,, 2024; Gordon et al.,, 2021; Maselli et al.,, 2023).
@@ 2.2 Experimental task and procedure
Participants could either deal with experimental conditions were all four areas are rich (referred to as the “rich environment” condition) or experiencing a combination of two rich and two poor areas (referred to as the “mixed environment” condition). Note that in the “mixed environment” conditions poor and rich areas are randomly distributed across the four areas. The distinction between rich and poor areas did not lie in the number of boxes contained in it, which is always five, but in the number of coins within them. In rich areas the amount of coins per box decreased linearly from 8 to 4 coins among the five boxes (i.e., the first opened box always delivered 8 coins, the second opened box 7 coins, etc.). Rather, in poor areas, the number of coins per box decreased linearly from 5 to 1 coin. Participants were informed about the presence of a spaceship in the central zone, which worked as a sensor (Figure 1.b; sensor not shown in the figure). The sensor provided participants with free information about the overall richness of the environment in the current trial, through colored (yellow) bars. Specifically, four long bars indicated the presence of a rich environment, in which all four areas are rich, whereas two long bars and two short bars indicated a mixed environment, in which two areas are rich and two are poor. Since all participants consulted the sensor at the beginning of each trial, we assume that they always knew whether the environment was currently “rich” or “mixed”. Note however that the sensor did not provide information about the specific areas designated as rich or poor for each trial – and rich and poor areas were allocated randomly at each trial. The sole method to ascertain the richness of an area involved actively exploring it and collecting boxes, which subsequently displayed the coin amounts contained in it (Figure 1.c).
@@ 4 Discussion
Human and other animals forage for valuable resources, such as food or other goods, in conditions in which the availability of the resources themselves and of foraging time might be limited. However, we still have a limited understanding of the ways the availability of resources and time influence human foraging strategies and whether participants are able to flexibly adapt their strategies to these different conditions – and whether they approximate optimal solutions provided by foraging theory (Charnov,, 1976).
To address these challenges, we designed a novel human foraging experiment that manipulates orthogonally the resource distribution and the time avilability. Concerning resources distribution, the experiment comprised trials in which all the areas were rich (rich environment) and trials in which two areas were rich and two were poor (mixed environment). Despite a fixed number of “treasure boxes” per area in all the trials (i.e., five boxes), the number of coins varied across rich areas (in which the number of coins per box decreased linearly from 8 to 4 coins) and poor areas (in which the number of coins per box decreased linearly from 5 to 1 coin). Concerning time availability, the experiment comprised two blocks, in which participants had longer (50 seconds) or shorter (25 seconds) foraging time, as signaled by a battery that was always visible to participants. To increase the realism of the situation, we adopted a video-game like scenario that incorporates some aspects of animal foraging, requiring participants not only to make abstract choices between areas and boxes, but also (virtually) navigating between areas, locating boxes and learning their possible locations.
@@
In sum, our results highlight the great flexibility of human foraging in the face of varying resource distributions, limited foraging time and learning opportunities. Furthermore, the results showcase the importance of studying foraging in rich environments that engage various cognitive processes, such as orientation and spatial navigation – and which afford learning and strategy adaptation over time.
@@ Acknowledgements
This research received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreements No. 952215 (TAILOR) to G.P.; the European Research Council under the Grant Agreement No. 820213 (ThinkAhead) to G.P.; the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, “EBRAINS-Italy”; Project PE0000013, CUP B53C22003630006, “FAIR”; Project PE0000006, CUP J33C22002970002 “MNESYS”) to G.P., PRIN PNRR P20224FESY to G.P., and the National Research Council, project iForagers. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.

### v3

@@ 1 Introduction
Foraging is a nearly universal behaviour in the animal kingdom and essential for survival of most species (Mobbs et al.,, 2018; Hayden,, 2018; Hills,, 2006; Calhoun and Hayden,, 2015; Stephens and Krebs,, 1986). Foraging problems encompass not just the search for survival-related resources (e.g., foraging for food) but also the search for information in the environment (e.g. in internet search (Pirolli,, 2007)) or in memory (e.g., “mental search” (Hills et al.,, 2015; Wilke et al.,, 2009; Todd and Hills,, 2020)).
In natural foraging problems, such as when animals navigate the environment to search for food (or other valuable resources), choice offers are presented serially and often found in clumps or patches, which gradually exhaust. This makes foraging scenarios conceptually different from classical economic tasks used to study value-based decisions, in which choice offers are typically presented simultaneously. Given their peculiarities, foraging choices are often conceptualized in terms of stay or leave decisions; namely, deciding whether to accept the available alternative (e.g., continue foraging in the same patch, in so-called patch-leaving problems) or reject it, in the hope that a better one would be found (e.g., move to another patch) (Charnov,, 1976; Hayden,, 2018; Hayden and Moreno-Bote,, 2018; Genovesio et al.,, 2014; Byrne and Whiten,, 1988). Stay-or-leave problems are extensively studied, both in humans and animals (e.g., mammals (Blanchard and Hayden,, 2015; Kane et al.,, 2022; Hayden et al.,, 2011; Gancarz et al.,, 2023; Shahidi et al.,, 2024), birds (Stephens and Dunlap,, 2009; Chandel et al.,, 2021)). Similarly, the trade-offs between exploration and exploitation that arise during foraging and related tasks have been widely studied in humans and other animals (Cohen et al.,, 2007; Mehlhorn et al.,, 2015; Addicott et al.,, 2017; Daw et al.,, 2006; Verschure et al.,, 2014; Maffei et al.,, 2015; Parr et al.,, 2022; Pezzulo et al.,, 2018; Friston et al.,, 2017)).
A formal solution to patch-leaving and other foraging problems is the Marginal Value Theorem of optimal foraging theory (Charnov,, 1976), according to which the most profitable moment to leave a patch (“giving-up time”) is when the reward in the current patch is less than the average reward rate of the environment (Constantino and Daw,, 2015). Neurally, the decision to stay or to leave is influenced by the time-linked reward representations at the level of the dACC (Wittmann et al.,, 2016). Furthermore, a recent PET study has demonstrated the role of dopamine, essential for motivational states that drive human decision-making dynamics (Le Heron et al.,, 2019), in regulating behavioral adaptations to changes in the reward environment by weighing reward benefits against temporal costs (Ianni et al.,, 2023).
@@
A series of human studies tested how environmental constraints – the most important being resource distribution – affect individuals’ patch-leaving strategies. Using a multiple-target visual search task, Cain and colleagues (Cain et al.,, 2012) demonstrated that foragers adapt to environmental statistics in a way that is consistent with, although not identical to, the optimal formulation: participants search more when expecting a greater amount of targets and modify their expectations in real time by considering the target distributions across trials. In a different experiment, a virtual patch-foraging task was used to assess the influence of environmental richness on foraging decisions (Constantino and Daw,, 2015). Results have shown that although participant’s behavior was in agreement with the optimal policy there was an “overharvesting” tendency, namely, a tendency to harvest a patch longer than it would be needed for reward maximization; see also (Hutchinson et al.,, 2008). Overharvesting has been explained as a strategy for reducing the cost of behavioral variability (Cash-Padgett and Hayden,, 2020), as a result of optimal learning and adaptability (Harhen and Bornstein,, 2023), or as a phenomenon occurring as a consequence of temporal cognition (Kendall and Wikenheiser,, 2022). Indeed, time definitely matters when making patch-leaving decisions; as the MVT states, longer travel times lead to longer patch residence times, which lowers the forager’s opportunity cost of time as well as the environment’s overall rate of reward (Bustamante et al.,, 2023).
@@
Like resources are limited in the environment, time is also a finite resource. Animals allocate different time foraging, depending on external constraints and they flexibly adjust their foraging behavior to time available (Krebs,, 1980; Swennen et al.,, 1989; Ydenberg et al.,, 1994; Ydenberg and Hurd,, 1998). For example, they tend to work harder when time available for foraging is more limited (Lewis et al.,, 2004). Time pressure also implies that decisions have to be made faster. A recent study employed a time bandit task to examine the effects of time pressure on exploration and decision making, found that in limited time conditions, participants earned less reward, were less sensitive to reward values in repeat choice behavior, were less likely to choose options that were associated with higher uncertainty, and explored less in uncertain scenarios. These findings can be explained by considering that under time pressure, participants tend to select simpler, lower-cost policies (Wu et al.,, 2022). In sum, the above studies contributed to shed light on how two environmental constraints – namely, resource distribution and time availability – affect human patch-leaving strategies. However, these two constraints have been manipulated in isolation, leaving open the question of how they might interact and whether participants are sensitive to both factors.
@@
Finally, most human foraging studies used simple paradigms, in which rewards are paired to buttons that are simultaneously present on the screen. It is unclear how well their results generalize to foraging scenarios that resemble more closely animal patch-living problems, in which rewards are presented serially, visibility is partial and which engage sophisticated cognitive processes, such as those required to learn and remember reward locations, orienting and navigating in space, etc. (Mobbs et al.,, 2018; Hayden,, 2018; Yoo et al.,, 2021; Cisek and Pastor-Bernier,, 2014). Crucially, these more challenging and ecologically valid scenarios present various sources of uncertainty about resource locations and distributions—for example, how many reward locations can be reached within limited time, where rewards are spatially located, and how to navigate efficiently to reach them—hence creating multiple occasions for learning. Therefore, they allow for novel questions about how foraging strategies adapt to environmental demands over the experiment, a factor often sidestepped in accounts like optimal foraging theory. For example, they enable researchers to study whether participants improve their foraging strategy and performance over time as they reduce uncertainty about key task dimensions, and whether, at some point, they match (or closely approximate) optimal solutions from foraging theory (Allen et al.,, 2024; Gordon et al.,, 2021; Maselli et al.,, 2023).
@@ 2.2 Experimental task and procedure
Participants could either deal with experimental conditions were all four areas are rich (referred to as the “rich environment” condition) or experiencing a combination of two rich and two poor areas (referred to as the “mixed environment” condition). Note that in the “mixed environment” conditions poor and rich areas are randomly distributed across the four areas. The distinction between rich and poor areas did not lie in the number of boxes contained in it, which is always five, but in the number of coins within them. In rich areas the amount of coins per box decreased linearly from 8 to 4 coins among the five boxes (i.e., the first opened box always delivered 8 coins, the second opened box 7 coins, etc.). Rather, in poor areas, the number of coins per box decreased linearly from 5 to 1 coin (i.e., the first opened box always delivered 5 coins, the second opened box 4 coins, etc.). Participants were informed about the presence of a spaceship in the central zone, which worked as a sensor (Figure 1.b; sensor not shown in the figure). The sensor provided participants with free information about the overall richness of the environment in the current trial, through colored (yellow) bars. Specifically, four long bars indicated the presence of a rich environment, in which all four areas are rich, whereas two long bars and two short bars indicated a mixed environment, in which two areas are rich and two are poor. Since all participants consulted the sensor at the beginning of each trial, we assume that they always knew whether the environment was currently “rich” or “mixed”. Note however that the sensor did not provide information about the specific areas designated as rich or poor for each trial – and rich and poor areas were allocated randomly at each trial. The sole method to ascertain the richness of an area involved actively exploring it and collecting boxes, which subsequently displayed the coin amounts contained in it (Figure 1.c).
@@ 4 Discussion
Human and other animals forage for valuable resources, such as food or other goods, in conditions in which the availability of the resources themselves and of foraging time might be limited. However, we still have a limited understanding of the ways the availability of resources and time influence human foraging strategies and whether participants are able to flexibly adapt their strategies to these different conditions as they learn more about the task and its statistics – and whether they approximate optimal solutions provided by foraging theory (Charnov,, 1976).
To address these challenges, we designed a novel human foraging experiment that manipulates orthogonally the resource distribution and the time availability. Concerning resources distribution, the experiment comprised trials in which all the areas were rich (rich environment) and trials in which two areas were rich and two were poor (mixed environment). Despite a fixed number of “treasure boxes” per area in all the trials (i.e., five boxes), the number of coins varied across rich areas (in which the number of coins per box decreased linearly from 8 to 4 coins) and poor areas (in which the number of coins per box decreased linearly from 5 to 1 coin). Concerning time availability, the experiment comprised two blocks, in which participants had longer (50 seconds) or shorter (25 seconds) foraging time, as signaled by a battery that was always visible to participants. To increase the realism of the situation, we adopted a video-game like scenario that incorporates some aspects of animal foraging, requiring participants not only to make abstract choices between areas and boxes, but also (virtually) navigating between areas, locating boxes and learning their possible locations. This scenario incorporates various sources of uncertainty, about (for example) resources distribution and time required to travel across boxes and areas, and hence various occasions for learning during during the experiment.
@@
One limitation of the current study is that we examined foraging decisions using a relatively simple model. This approach was taken because we aimed to investigate the extent to which people approximate optimal foraging strategies during the experiment, assuming no uncertainty except for the location of rich and poor areas. Future studies could formalize alternative strategies that individuals might adopt when deciding how to forage and when to leave a patch. Previous research has identified possible foraging rules (e.g., giving-up time and area-restricted search in (Wilke et al.,, 2009)) that could be adapted to our task and used for model comparison, providing further insight into how people forage and whether their strategies change between early and later phases of foraging. Additionally, incorporating learning models could offer a more detailed examination of explore-exploit dynamics during foraging (Friston et al.,, 2016; Sutton,, 2018). Finally, future studies might adapt our setup to clinical or subclinical populations to investigate under which conditions they develop adaptive or maladaptive foraging strategies (Addicott et al.,, 2017; Barack et al.,, 2024; Pellicano et al.,, 2011).
In sum, our results highlight the great flexibility of human foraging in response to varying resource distributions, limited foraging time, and learning opportunities. Furthermore, they showcase the importance of studying foraging in rich environments that engage various cognitive processes, such as orientation and spatial navigation, while also affording learning and strategy adaptation over time. These ecologically valid contexts enable new questions about how foraging strategies adapt to the physical properties of the environment, revealing subtle aspects often overlooked in accounts like optimal foraging theory. For example, our experiment suggests that people transition from an initial exploration phase—where they reduce uncertainty about key aspects of the task, such as resource locations and travel times—to a subsequent exploitation phase, in which they leverage their knowledge to deploy more effective foraging strategies. These learning and adaptation effects in foraging present a valuable avenue for future research.
@@ Acknowledgements
This research received funding from the European Union’s Horizon 2020 Framework Programme for Research and Innovation under the Specific Grant Agreements No. 952215 (TAILOR) to G.P.; the European Research Council under the Grant Agreement No. 820213 (ThinkAhead) to G.P.; the Italian National Recovery and Resilience Plan (NRRP), M4C2, funded by the European Union – NextGenerationEU (Project IR0000011, CUP B51E22000150006, “EBRAINS-Italy”; Project PE0000013, CUP B53C22003630006, “FAIR”; Project PE0000006, CUP J33C22002970002 “MNESYS”) to G.P., PRIN PNRR P20224FESY to G.P., and the National Research Council, project iForagers. The funders had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript. We used a Generative AI model to correct typographical errors and edit language for clarity.